mathslover (mathslover):

Hi friends another tutorial on solving linear equations in 1 variable

mathslover (mathslover):

i have seen many of the students asking questions like this \[\huge{x+3=9x}\] what is x ? so this is a tutorial for those students

mathslover (mathslover):

i will make this tutorial easy for them by just giving the explanations in points LINEAR EQUATIONS IN 1 VARIABLE This topic is much important for beginners in algebra. .. it contains some methods or processes that should be remembered : 1) what is equation ? :: equation word is derived from the word equate ... if a expression contains equal to sign ( = ) than it can be refferred to as equation for example \(\huge{x+3=2 , x+2 = 4}\) these are the quations 2) LHS or RHS : Left hand side and Right Hand Side are the two important terms that one should remember .. left hand side means the expression in left side of the equation and Right Hand Side means the right part of the equation

mathslover (mathslover):